This one is an application problem: 16
During one month, a homeowner used 500 units of electricity and 100 units of gas for a total cost of $352. The next month, 400 units of electricity and 150 units of gas were used for a total cost of $304. Find the cost per unit of gas.
I am really unsure how to go about solving this

Answer by richwmiller(17219)

500e+100g=352
400e+150g=304
g gas = 8/25=$.32 e electricity = 16/25=$.64